# A Topic-wise Exploration of the Telegram Group-verse

### v1

@@ 4.2.1. Telegram Bot usage and user activity
Figure 1. ECCDF of the fraction of messages sent by bots in a group.
Telegram group administrators can incorporate bots into their groups. Bots offer a wide array of functionalities, from welcoming newcomers with group rules to responding to user commands, from collecting statistics to moderating messages. Telegram bots are in fact quite popular: only 10.2% of groups in our dataset do not include bots; in the median, there are four bots per group. Interestingly, 1.5% of groups have 20 bots (the maximum allowed by Telegram). Some bots enjoy significant popularity: Combot(com, 2024) and MissRose_bot(mis, 2024) are installed in 145 and in 129 groups, respectively. Both provide moderation services, analytics, and anti-spam features.
Bots footprint is not negligible: in groups with bots, they generate on average 8.6% of messages, with notable variations across topics. Figure 1 shows, for 4 topics, the Empirical Complementary Cumulative Distribution Function (ECCDF) of the fraction of messages sent by bots for different groups. In Linguistics, bots generate the highest fraction of messages (brown dotted curve). In fact, some bots are integral to learning platforms and merit examinations. For instance Quizbot(qui, 2024) is widely deployed in Linguistic and Education groups (27.5% and 28.6%, respectively). It generates an average of 39.0% and 13.0% of messages. In one Linguistics group it generates 91% of messages. Conversely, Politics groups see the smallest fraction of messages generated by bots (green dotted line), possibly testifying to a higher user engagement [^5] For messages sent by a user account, we cannot distinguish between messages sent by a human or by an automated system. in political groups than in other topics. Curiously, there is a quite large fraction of groups with bots that simply collect statistics or moderate the group without sending any message (leftmost part of Figure 1). For the remainder of our analysis, we ignore messages sent by Telegram bots.
Focusing on the amount of messages actual users generate, we observe that there are few users sending thousands of messages, while the majority are not active (see column 4 in Table 4.1.1) or send few messages. Indeed, the Empirical Probability Density Function (EPDF) of the number of messages generated by users follows a heavy-tailed shape that can be fit by a Pareto distribution with $\alpha=1.9$. Remarkably, the fittings of the per-topic EPDFs are very similar, hinting at the universality of these behaviours.
@@ 4.2.2. Language
The next question we answer is what are the languages people speak in each topic and group. For each textual message in a group, we associate a language by employing the FastText language identification library (Joulin et al., 2017) to obtain the distribution of languages for each group. In Figure 2 we report the breakdown of the most popular language for groups on the same topic:
$\bullet$ English (a global language) and Russian (being Telegram very popular in Russia) are the two most popular languages. Their share changes based on the topic. For instance, most groups in Bookmaking and Darknet have Russian messages. Conversely, the majority of Education and Cryptocurrency groups are in English, possibly due to the worldwide interest in such topics.
$\bullet$ Despite Telegram’s restricted use in Iran, some popular groups are in Persian, especially in Politics. This is in line with the claim that Telegram is among the platforms to evade state censorship in Iran (Akbari and Gabdulhakov, 2019).
$\bullet$ Telegram is blocked in China. In fact, it does not appear to be popular for discussing topics in Chinese. Yet, few groups have Chinese as the dominant language.
We observe that in almost half of the groups, more than 75% of the messages are written in the same dominant language. Still, in all groups and topics, we observe messages written in other languages, hinting to a global user population. Linguistics is the topic where groups contain the largest mix of languages, which supports the anecdotal observation of users practising foreign languages and mixing messages in their native language in such groups. In contrast, both Darknet and Technology stand out with more than 58% of the groups having more than 80% of the messages in a single language (English). This agrees with the intuition that technology-related discussions are carried over in English.
@@ 4.2.3. Message length
Figure 3 shows the ECCDF of the length of textual messages for each topic. Since message length may be influenced by language, we consider only messages written in English-dominated groups. We observe some great distinctions across topics. On one hand, Darknet groups are dominated by very long messages (80% longer than 100 characters). A manual check unveils that most messages contain samples of illicit content people trade. The same can be said for Bookmaking and Technologies groups, where long messages describe bookmaking websites, experiences and results of betting, or devices to sell. In Erotic, people advertise their services. Conversely, Linguistics and Politics groups are dominated by very short messages in which people debate or provide suggestions (80% shorter than 30 characters). In a nutshell, the presence of “advertisement” messages inflates their length, while discussion-driven groups see a predominance of short messages.
At last, steps in the distribution signal the presence of repeated automated messages (mostly spam).
@@ 4.2.4. Usage of non-textual elements
We now broaden our analysis to consider non-textual elements. Specifically, we extract, for each group, the fraction of messages containing images, external links, voice messages, polls, GIFs, stickers, videos, and emojis. How are these elements used? To gauge this, we compute the average usage fraction over all groups of a given topic (macro average). Figure 4 visually compares two pairs of selected topics using radar charts. Table 3 and Figure 8 in Appendix B provides the complete set of results. Some interesting findings emerge:
$\bullet$ Politics groups represent the typical average usage of non-textual elements: 20–30% of messages contains emojis; 10% of messages shares an image; stickers are more popular than GIFs; few messages contain voice content; polls are mostly an unused feature (present only in Linguistics groups).
$\bullet$ Cryptocurrencies groups represent some mixed usage: no videos and voice messages, fewer photos and emojis but more stickers and GIFs.
$\bullet$ Groups in Video and Films and Politics have very similar usage patterns (i.e. radar shape), though, surprisingly, the former has fewer videos (present in only 1% of the messages) — see Section 5.
$\bullet$ Erotic groups present the minimum usage of non-textual elements: no stickers, no GIFs, while photos are found in 6% of messages. Surprisingly, we see very few links to external platforms. Indeed, readers are invited to contact “advertisers” via private chat.
Notice that sending stickers requires manual actions hardly automatizable. Also, stickers are commonly used as reactions to other messages. Their prominent use in a group or topic may testify to a larger fraction of messages being sent by real users, or to a more confidential exchange: Erotic, Darknet and Technologies have the least fraction of stickers and are dominated by ad-style messages. Conversely, in Bookmaking and Cryptocurrencies people use more stickers as reactions for suggestions.
@@ 5.1. Video size and duration
We focus our analysis on groups in the three topics with the largest share of videos: Politics, Video and Films, and Erotic. Although Figure 4 suggests some similarities in the amount of shared videos, a closer examination of the video duration and file size reveals noteworthy differences in goals and types of shared videos.
The left plot of Figure 5 compares the video duration. Observe how videos shared in Video and Films are notably longer than in other topics, with peaks roughly around the 60- and 120-minute marks (notice the log-y scale). Manual inspection confirms that people share entire TV series episodes and movies. Their total volume amounts to 7.45 TB of data.
A comparison of video duration with file size — see the right graph of Figure 5 — exposes an noticeable difference between topics: videos shared in Erotic groups have significantly higher video rate (thus quality), as evidenced by the steeper slope of the regression line.
@@ 5.2. External URL lookup
Analyzing the sharing of links to external websites (i.e., different from telegram.me and t.me) allows us to understand how Telegram users may be redirecting (or driving) attention (and traffic) to other websites.
@@
In Figure 6, we present the average frequency at which a domain appears in a given topic. We focus on the union of the 5 most popular domains across each topic. These collect from $\approx$20to $\approx$50of links — with a fragmented list of other platforms (bottom row). The three most frequent platform are social networks: X, YouTube and Instagram. Usage varies a lot: X sees significant use in Cryptocurrency; YouTube and, to a lesser extent, Instagram are the two most transversal platforms. Conversely, in line with their discussion topics, Technology and Software and Application show a significant usage of GitHub pages. The telegra.ph open and anonymous publishing platform is used in Software & Applications, e.g. to share installation guides and tutorials.
@@ 6. Conclusion
In this paper, we offered a first-of-its-kind transversal analysis of Telegram as observed through the lens of public groups retrieved from TGstat. We observed a large diversity of the usage that communities do of the platform capabilities, which in turn reflects the type of content and goals of sharing: for instance, bots may have a huge footprint — with up to 80–90% of bot-generated messages. User messages length varies a lot — with the self-advertisement-driven content inflating message length; video sharing is quite common — in Politics these are short, while in Video and films entire movies get shared — Erotic having the highest quality.
All in all, our work, albeit preliminary, paves the road to the investigations and understanding of how people use the Telegram platform. We hope our work can foster discussion towards a deeper understanding of this platform.
Appendix A Ethics
In our work, we take ethics under utmost consideration throughout the whole process.

### v2

@@ 4.2.1. Telegram Bot usage and user activity
Figure 1. Fraction of messages sent by bots in a group across topics.
Telegram group administrators can incorporate bots [^5] Note that we use the term bot to refer exclusively to official bots and not to users exhibiting bot-like automated behaviour. into their groups, offering a wide set of functionalities, from welcoming newcomers with group rules to responding to user commands, from collecting statistics to moderating join requests and messages. Telegram bots are quite popular: only 10.2% of groups in our dataset do not include bots; in the median, there are four bots per group. Interestingly, 1.5% of groups have 20 bots (the maximum allowed by Telegram). Some bots enjoy significant popularity: Combot (com, 2024) and MissRose_bot (mis, 2024) are present in 145 and 129 groups, respectively. Both provide moderation services, analytics, and anti-spam features.
Bots footprint is not negligible: they generate on average 8.6% of messages, with notable variations across topics. Figure 1 shows, for 4 topics, the Empirical Complementary Cumulative Distribution Function (ECCDF) of the fraction of messages sent by bots for different groups. The largest fractions of messages sent by bots are in Linguistics groups. Some of these bots are integral to learning platforms and merit examinations. For instance Quizbot (qui, 2024) is widely deployed in Linguistics and Education groups (27.5% and 28.6%, respectively). It generates 39.0% and 13.0% of messages. In one Linguistics group it generates 91% of messages. Conversely, Politics groups see the smallest fraction of messages generated by bots (green dotted line), possibly testifying to a higher user engagement [^6] For messages sent by a user account, we cannot distinguish between messages sent by a human or by an automated system. in political groups than in other topics. Curiously, there is a quite large fraction of groups with bots that simply collect statistics or moderate the group without sending any message (leftmost part of Figure 1). For the remainder of our analysis, we remove messages sent by Telegram bots.
Focusing on the number of messages actual users generate, we observe that there are a few users sending thousands of messages, while the majority are not active (see column 4 in Table 4.1.1) or send few messages. Indeed, the Empirical Probability Density Function (EPDF) of the number of messages generated by users follows a heavy-tailed shape that can be fit by a Pareto distribution with $\alpha=1.9$. Remarkably, the fittings of the per-topic EPDFs are very similar, suggesting the universality of these behaviours, as widely acknowledged in the literature(Barabasi, 2005; Rybski et al., 2009).
Main takeaways: official bots’ impact on the activity of Telegram is significant (8.6% of messages), with some moderation bots being present on a wide variety of groups. Some topics are more prone to bots offering also specific features (e.g., QuizBot generating 39% of messages in Linguistics’ groups.
@@ 4.2.2. Language
The next question we answer is what are the languages people use in each topic and group. For each textual message in a group, we associate a language by employing the FastText language identification library (Joulin et al., 2017) to obtain the distribution of languages for each group. In Figure 2 we report the breakdown of the most popular language for groups on the same topic:
$\bullet$ English (a global language) and Russian (being Telegram very popular in Russia) are the two most popular languages. Their share changes based on the topic. For instance, most groups in Bookmaking and Darknet have a lot of Russian groups. Conversely, the majority of Education and Cryptocurrency groups are in English, possibly due to the worldwide interest in such topics.
$\bullet$ Despite Telegram’s restricted use in Iran, some popular groups are in Persian, mostly in Politics. This is in line with the claim that Telegram is used to evade state censorship in Iran (Akbari and Gabdulhakov, 2019).
$\bullet$ Although Telegram is blocked in China (tel, 2024a), we do find groups with Chinese as the dominant language (up to 21% of Erotic groups).
We observe that in almost half of the groups, more than 75% of the messages are written in the same dominant language. Still, we observe messages written in other languages, hinting at a global user population. Linguistics is the topic where groups contain the largest mix of languages, which supports the anecdotal observation of users practising foreign languages and mixing messages in their native language in such groups. In contrast, both Darknet and Technology stand out with more than 58% of the groups having more than 80% of the messages in a single language.
Main takeaways: English and Russian are the most widely used languages, dominant in 34.1% and 38.4% of groups, respectively. Notably, Persian (3.9%) and Chinese (6.0%) also appear frequently, despite Telegram restrictions in Iran and China. Many groups mix multiple languages within their discussions.
@@ 4.2.3. Message length
Figure 3 shows the ECCDF of the length of textual messages for each topic. Since message length may be influenced by language, we consider only messages written in English-dominated groups. We observe some great distinctions across topics. On one hand, Darknet groups are dominated by very long messages (80% longer than 100 characters). A manual check unveils that most messages contain samples of illicit content people trade, such as bank account credentials, credit cards, passports and SSNs. Conversely, in Bookmaking and Technologies groups, long messages describe bookmaking websites, betting experiences and results, or offer details about devices for sale. In Erotic groups, long messages are used to advertise services and sell content, some of them repeated multiple times (observe the steps in the ECCDF). On the other hand, Linguistics and Politics groups are dominated by very short messages in which people debate or provide quick suggestions (80% shorter than 30 characters). At last, steps in the distribution might suggest the presence of repeated automated messages (mostly spam), further analysed in Section 5.3.
Main takeaways: There appears to be a correlation between message length and promotional content, with longer messages linked to advertisements, while discussion-driven groups to shorter messages (e.g., compare Bookmaking and Linguistics in Figure 3).
@@ 4.2.4. Usage of non-textual elements
We now broaden our analysis to consider non-textual elements. Specifically, we extract, for each group, the fraction of messages containing images, external links, voice messages, polls, GIFs, stickers, videos, and emojis. How are these elements used? To gauge this, we compute the fraction of messages with such element in a group and compute the average over all groups of a given topic (macro average). Figure 4 visually compares two pairs of selected topics using radar charts. Table 2 and Figure 4.1.1 in Appendix A provides the complete set of results. Some interesting findings emerge:
$\bullet$ Politics groups represent the typical usage of non-textual elements: 20–30% of messages contain emojis; 10% of messages share an image; stickers are more popular than GIFs; few messages contain voice content; polls are mostly an unused feature (polls are in fact only used in Linguistics groups).
$\bullet$ Cryptocurrencies groups represent some mixed usage: no videos and voice messages, fewer photos and emojis but more stickers and GIFs than average.
$\bullet$ Groups in Video and Films and Politics have very similar usage patterns (i.e. radar shape), though, surprisingly, the former has fewer videos (which are present in only 1% of the messages) — see Section 5.
$\bullet$ Erotic groups present the minimum usage of non-textual elements: no stickers, no GIFs, while photos are found in 6% of messages. Surprisingly, we see very few links to external platforms. In fact, readers are invited to contact advertisers via private chat.
Notice that sending stickers requires manual actions hardly automatizable. Also, stickers are commonly used as reactions to other messages. Their prominent use in a group or topic may testify to a larger fraction of messages being sent by real users, or to a more confidential exchange: Erotic, Darknet and Technologies have the least fraction of stickers and are dominated by ad-style messages. Conversely, in Bookmaking and Cryptocurrencies people use more stickers as reactions for suggestions. This argument was highlighted on other platforms such as WhatsApp where, for instance, in political groups, stickers are rarely forwarded, suggesting that users often save and maintain their collection of preferred stickers for future use rather than relying on the platform’s sharing tools (Melo et al., 2024), further highlighting a more human-based interaction pattern.
Main takeaways: the media-sharing patterns across topics show big differences based on the topic under observation. No common habits emerge.
@@
Table 2. Per-topic detailed metrics of usage of non-textual elements in messages.
Topic Poll (%) Voice (%) Image (%) Emoji (%) Video (%) Sticker (%) GIF (%) Link (%)
Education 0.01 0.02 5.13 12.43 0.04 0.01 $<$0.01 0.06
Bookmaking $<$0.01 0.03 8.67 19.81 0.40 0.72 0.38 0.04
Crypto $<$0.01 $<$0.01 3.89 12.54 0.04 0.81 0.39 0.14
Technologies $<$0.01 $<$0.01 6.12 8.22 0.10 0.09 0.05 0.29
Darknet $<$0.01 $<$0.01 5.44 56.51 0.34 $<$0.01 0.03 0.02
Software $<$0.01 $<$0.01 6.50 9.61 0.42 0.42 0.10 0.36
Video&Films $<$0.01 0.02 7.61 15.10 1.15 0.58 0.17 0.10
Politics $<$0.01 0.07 9.69 22.17 3.03 0.25 0.17 0.26
Erotic $<$0.01 $<$0.01 6.05 17.29 0.80 0.03 0.02 $<$0.01
Linguistics 0.10 0.08 2.34 15.34 0.33 0.16 $<$0.01 0.04
All Topics $<$0.01 $<$0.01 6.02 14.65 0.23 0.23 0.09 0.07
@@ 5.1. Video size and duration
We focus our analysis on the three topics with the largest share of videos: Politics, Video and Films, and Erotic. Although Figure 4 suggests some similarities in the amount of videos people directly share on Telegram, a closer examination of the video duration and file size reveals noteworthy differences in goals and types of shared videos, again showcasing notable differences in users’ behavioural patterns among different topics.
The left plot of Figure 5 compares the video duration, by showing the fraction of videos by length. Observe how videos shared in Video and Films are notably longer than in other topics, with peaks roughly around the 60- and 120-minute marks (notice the log-y scale). Manual inspection confirms that people share entire TV series episodes and movies. Their total volume amounts to 7.45 TB of data. Erotic and Politics tend to have much shorter videos.
Compare now this video duration with its file size — see the right graph of Figure 5, a scatter plot with regression lines, where each point represents a shared video. Again, we observe a noticeable difference between topics: videos shared in Erotic groups have significantly higher video rate (thus quality), despite shorter durations, as evidenced by the steeper slope of the regression line. In contrast, videos shared in Politics groups tend to be shorter, with limited attention given to production quality. Their primary function is to serve immediate needs, such as delivering brief announcements, political speeches, or viral clips intended to shape opinions. A sample of these videos reveals content related to the Russian-Ukrainian conflict, speeches by politicians, and clips with political undertones designed to provoke indignation.
Main takeaways: video quality (Figure 5, left) and duration (Figure 5, right) vary based on the goal of sharing such a video, from entire movies to a few seconds clips to shape the debate.
@@ 5.2. Links to external domains
We now extend the analysis to the sharing of links to external websites (i.e., different from telegram.me and t.me) to assess how Telegram users wish to redirect (or drive) attention to other websites, and the interplay with external Web resources .
@@
In Figure 6, we present the average frequency at which a domain appears in a given topic. We focus on the union of the 5 most popular domains across each topic. These cover from $\approx$20to $\approx$50of links — with a heavy-tailed presence of other platforms (bottom row). The three most frequent platforms are social networks: X, YouTube and Instagram. Usage varies a lot: X sees significant use in Cryptocurrency groups; YouTube and, to a lesser extent, Instagram are the two most transversal platforms. Conversely, in line with their discussion topics, Technology and Software and Application groups show a significant usage of GitHub pages. The telegra.ph, an open and anonymous publishing platform, is very often used in Software & Applications groups, e.g., to share anonymously installation guides and tutorials. All in all, we observe a very diverse sharing of information from extensive platforms, with each topic having different preferential means to refer to external outlets.
Finally, we expand bit.ly URLs, finding that only 4.2% pointed to popular domains (i.e., those in the union of the five most frequent domains per topic), while 67.4% led to less common websites with few occurrences. Additionally, 28.4% of these URLs directed to expired domains. As bit.ly URLs represent a small fraction of the total, their expansion minimally impacts domain rankings but rather highlights the use of URL shorteners as a noteworthy behavior.
Main takeaways: YouTube consistently serves as the most prevalent domain across all topics to share external content (largest share of links across each but one topic). A notable exception is the cryptocurrency discussions, where X takes precedence (11.9% vs. 6%). Some platforms are popular in some topics, with again few common habits (except YouTube, see Figure 6).
5.3. Repeated sharing behavior
Figure 7. Cumulative time series of unique URLs appearing at least twice in the dataset, along with all their subsequent repetitions. The zoom highlights the first two days.
Delving into the analysis of the links to external domains, we now look at possible patterns from the repeated sharing of the same URL. In total, there are around 239,000 unique URLs in our dataset, around 43,000 (18%) of which are shared multiple times during the period of analysis. We here focus on these duplicated URLs.
In Figure 7 we report a scatter plot where every dot represents the sharing of a given URL (y-axis) at a specific timestamp (x-axis). We sort the unique URLs by their first appearances in our data. We distinguish if the URL is shared multiple times by the same sender (old sender) or a new one (new sender), and if the URL is shared multiple times in the same group (old group) or in a new one (new group). A sender is new when they first send a given URL to a given group. If they send the same URL to the same group again, they become old. If they send the same URL to another group for the first time, they are new again.
We find that the large majority (93.2%) of all duplicated URLs correspond to users posting the same URLs in groups they have already posted the URLs in. This hints at a spamming activity by regular accounts that programmatically share the same URL with the same audience, to maximise the reach of the content. Most of these repeated URLs appear in groups about Bookmaking (54%), Cryptocurrencies (21%) and Darknet (14%) topics. These groups are in fact full of ads and spam messages.
By manually analysing the most replicated URLs, we observe that many of them advertise Social Media Managing services, which promise to boost one’s interactions (followers, likes,etc.) through the use of likely fake accounts. Curiously, the most repeated URL appears more than 27,000 times in a single group over two months.
Our analysis reveals also suspicious patterns. The red line in the inset of Figure 7 shows an uninterrupted sequence of messages with the same URL sent within the same Bookmaking group by 1,498 distinct users. This suggests a coordinated network of likely fake users promoting a betting website. In the first week of April, we also observe a sudden burst of shared URLs that are shared for a few days. These URLs refer to Reddit comments or posts arguably sent by their author who asks for “upvotes” to other members of a Cryptocurrencies group, suggesting an attempt to manipulate content curation algorithms as shown in (Weerasinghe et al., 2020).
Figure 8. Cumulative time series of unique long messages (over 50 characters) repeated at least twice in the dataset, including all subsequent identical reposts.
After detecting spamming behaviour in URL sharing, we investigate whether similar patterns occur with textual messages. To filter out trivial repetitions (e.g., ”ok” or ”good morning”), we focuse on messages longer than 50 characters and analyse perfect duplicates with identical content. As shown in Figure 8, duplication is common even for long messages, with nearly 500,000 messages repeated at least once across 97.8% of the groups, accounting for 65% of all long messages. Unlike URLs, duplicated messages are more evenly spread across topics, with the Darknet topic showing the highest repetition (19%) and Linguistics, Video and Films, and Education having the fewest (under 1%).
Although Telegram allows message forwarding, only 13.7% of these repeated messages are forwarded; most are posted directly by the same users. Similar to URL patterns, most duplicated messages come from users repeatedly posting the same content in the same group, further indicating spamming behavior.
Suddenly, starting from the last week of March 2024, a great portion of duplicated messages is reposted in the same group by users who had not previously shared them (new senders old group, red pattern). Interestingly, this phenomenon happens in a single group. This group (“kadyrov_95chat”) contributes to 4.3% of all the repeated messages and 84% of all the repeated messages sent by different users in the same group. Predominantly in Chechen, messages often include congratulations, generic wishes, and praise for the Chechen Republic and its government. We conjecture that this behaviour is caused by some sort of spam bot networks, which tend to copy-paste the same message multiple times. However, we acknowledge that we could not retrieve messages sent before March 27th, leaving it unclear whether the group was previously inactive or if earlier messages were deleted before our data collection began. This highlights the importance of tracking a group activity over time, to retrieve its historical messages.
Main takeaways: we observe indications of spamming behaviour in the sharing of both URLs (Figure 7) and textual messages (Figure 8). Most repeated content comes from the same single users (93.2% and 89.8% respectively). Yet, evidence of coordinated group of spammers clearly emerges (e.g., red line in the inset of Figure 7). This should encourage further research to characterise this behaviour and its impact on user participation and group dynamics.
5.4. Time Lag in YouTube Video Sharing
As shown in Figure 6, youtube.com is the most consistent domain shared on Telegram groups across topics. It accounts for the 8.1% of the unique links shared in our dataset. For each posted video link, we collect its publication timestamp on YouTube through the YouTube Data API v3.777https://developers.google.com/youtube/v3?hl=en Then, we measure the elapsed time to its first appearance in all of our Telegram groups. [^8] If appearing in multiple groups, we evaluate the difference multiple times. In Figure 9, we present the ECCDF for the topics with the most diverse distributions. Considering videos shared in Education groups, more than 40% of them are shared on Telegram less than 24 hours after their publication on YouTube. This suggests a community interested in fresh content. Most videos consist of classes and tutorials offering guidance on how to approach specific exams. The Video and films topic, in turn, exhibits a great interest in old videos: more than 30% of the videos have been on YouTube for almost three years (mostly old movies). This indicates that this topic is not recency-bound.
In the Technologies groups, we observe a limited interest in new videos, with only 20% of shared content appearing within 10 days of publication. This is surprising, as one would expect a high level of engagement from the community regarding the latest developments in the field. One explanation lies in the predominance of video tutorials on technology topics, which retain their relevance over time. In fact, most URLs refer to educational content or guides in the technology field, while some present consumer technology reviews.
Finally, it is noteworthy to focus on the knee of the ECCDF of videos shared in the Cryptocurrencies groups: here videos older than one year are less and less shared. This can be due to the variety of Cryptocurrencies and market-related suggestion videos, whose content naturally loses interest over time.
Main takeaways: the distribution of the elapsed time between publication on YouTube and sharing on Telegram (Figure 9) again showcases context-based patterns that depend on the communities interacting on the platform.
Figure 9. Elapsed time between a video publication on YouTube and its first appearance in Telegram groups, considering 4 topics (log x-scale).
@@ 6. Conclusion
In this paper, we conducted a comprehensive, topic-wise analysis of Telegram public groups retrieved from TGStat, providing a transversal view of platform usage across diverse discussion topics. Our work contributes to the understanding of how Telegram is used across various domains, shedding light on usage patterns that had not been explored in prior studies. Furthermore, we develop an open-source crawler and dataset that contributes valuable tools for future researchers to explore Telegram usage.
Our analysis reveals a significant heterogeneity in how different communities leverage the platform’s capabilities. Bots, for instance, play a crucial role in many groups, with some topics — such as Linguistics and Education—showing up to 80–90% of messages generated by bots. This highlights the importance of automated interaction in certain topics, where bots are integral to group functionality. Conversely, in Politics groups, bot activity is notably lower, suggesting higher levels of direct user engagement in discussions.
Media sharing varies significantly across topics. In Politics groups, videos are short and focused on quick messages or viral clips, while in Video and Films groups, entire movies and TV shows are shared, contributing to a larger data volume. Erotic groups prioritize high-resolution videos despite shorter durations. Non-textual elements like stickers, emojis, and links also differ, reflecting diverse community dynamics and interaction styles.
In terms of user activity, the length of user-generated messages differs greatly. In Darknet groups, messages are particularly long, often containing detailed information about illicit goods and services, while in discussion-driven groups like Politics and Linguistics more concise messages prevail. Moreover, we identified substantial repeated sharing behaviour, particularly in Bookmaking, Cryptocurrency, and Darknet groups, where spamming and coordinated actions (e.g., repeated posting of URLs or messages by different users) are common. We highlighted the presence of bot-like networks or coordinated manipulation efforts within specific communities.
Overall, our study offers a broad, multi-faceted exploration of Telegram group usage that contrasts sharply with previous studies, which were more limited in scope. The diversity of patterns we uncovered demonstrates how versatile and complex Telegram usage can be, depending on the topic at hand.
Future research directions include identifying and quantifying the key factors that drive user engagement and examining the processes of influence and information dissemination. Additionally, further analysis of content through NLP techniques could shed light on shifts in discussion subjects.
7. Ethical aspects
In our work, we take ethics under utmost consideration.